OpenStudy (anonymous):

find x,y intercepts..f(x)=-x^2+7x+30. I know y just need help with x

OpenStudy (mertsj):

Find the y intercept by letting x = 0. Find the x intercepts by letting y = 0.

OpenStudy (anonymous):

x^2+7x+30 = 0 Find two possible values of x.

OpenStudy (anonymous):

do i need to factor first? cuz if i do i dont know what to put for 30

OpenStudy (anonymous):

Use the formula.

OpenStudy (anonymous):

yea i gotta find two points for x

OpenStudy (mertsj):

f(x)=-(x^2-7x-30)=-(x-10)(x+3)=0 x=10, -3

OpenStudy (anonymous):

wait would it be 10 and 3?

OpenStudy (anonymous):

No, that gives -30 not +30. Unless the problem is wrong.

OpenStudy (mertsj):

No. 10 and -3

OpenStudy (anonymous):

oh yes -3

OpenStudy (mertsj):

Factor out the negative 1 as I showed and then factor x^2-7x-30

OpenStudy (anonymous):

so when i factor do i just not do the 7x?

OpenStudy (anonymous):

10 and -3 are NOT right if the problem is x^2+7x+30=0.

OpenStudy (mertsj):

You are factoring this trinomial: x^2-7x-30

OpenStudy (anonymous):

and the y intercept is 30?

OpenStudy (mertsj):

The problem is - x^2+7x+30

OpenStudy (anonymous):

wait if it was minus it would be 10,-3?

OpenStudy (anonymous):

Got it. I missed that. Thanks!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

ok i got confused for a minute

OpenStudy (mertsj):

All I know is that the problem you posted said: f(x)=-x^2+7x+30 and the answers to that problem are 10 and -3 and I have shown you how to do it.

OpenStudy (anonymous):

um yea thats the answer i got, thanks